# SVD-LLM: Truncation-aware Singular Value Decomposition for Large Language Model Compression

### 12 Mar 2024 v1

@@ 1 Introduction
We compare SVD-LLM with three SVD-based methods for LLM compression, including vanilla SVD as well as state-of-the-art methods FWSVD and ASVD. To demonstrate the generability of SVD-LLM, we conduct our evaluation on a total of $11$ datasets and seven models from three different LLM families (LLaMA, Vicuna, and Mistral) at four different scales (7B, 13B, 30B, and 65B). We highlight five of our findings: (1) SVD-LLM consistently outperforms vanilla SVD, FWSVD, and ASVD across all $11$ datasets, three different LLM families, and four different scales. (2) None of the existing SVD-based LLM compression methods report their performance on model compression ratios above 20%. SVD-LLM exhibits significant advantages over vanilla SVD, FWSVD, and ASVD under high compression ratios from 30% to 60%. (3) SVD-LLM also exhibits superiority over state-of-the-art in terms of compression speed. Specifically, when compressing LLaMA-7B under 20% compression ratio on an A100 GPU, ASVD takes about $5.5$ hours whereas SVD-LLM completes the compression process in $15$ minutes. (4) Moreover, SVD-LLM can benefit other LLM compression methods. Our evaluation results show that SVD-LLM is able to further enhance the compression performance of state-of-the-art quantization (GPTQ (Frantar et al., 2022)) and parameter pruning-based (LLM-Pruner (Ma et al., 2023)) LLM compression methods. (5) Lastly, SVD-LLM brings additional benefit beyond compressing the sizes of LLMs, and is also able to reduce the footprint of KV cache during inference at runtime.
@@ 3.1 Truncation-Aware Data Whitening
Key Design: The key idea of SVD-LLM is to incorporate a truncation-aware data whitening technique that ensures a direct mapping between singular values and compression loss. To achieve this, SVD-LLM change the whitened activation $\tilde{X}=S^{-1}X$ to be orthogonal so that each channel is independent of each other, i.e., $\tilde{X}\tilde{X}^{T}=S^{-1}XX^{T}(S^{-1})^{T}=I$. Though the matrix $S$ is not unique, we found that the Cholesky decomposition (Meyer, 2023) of $XX^{T}$ is a good solution due to its numerical stability. After that we perform SVD on $\tilde{W}=WS$ instead of the original weight $W$ to obtain three decomposed matrices $\tilde{U},\tilde{\Sigma},\tilde{V}$, where $\tilde{U}=[\tilde{u_{1}},\tilde{u_{2}},\tilde{u_{3}},...,\tilde{u_{r}}]$, $\tilde{\Sigma}=\text{diag}(\tilde{\sigma_{1}},\tilde{\sigma_{2}},\tilde{\sigma_{3}},\cdots,\tilde{\sigma_{r}})$, and $\tilde{V}=[\tilde{v_{1}},\tilde{v_{2}},\tilde{v_{3}},...,\tilde{v_{r}}]$. Lastly, the smallest singular values are truncated to compress $\tilde{\Sigma}$ to $\tilde{\Sigma}_{1}$ and obtain the compressed weight $\tilde{W^{\prime}}=\tilde{U^{\prime}}\tilde{\Sigma}_{1}\tilde{V}^{T}$.
@@ 4.1 Experimental Setups
Models and Datasets. To demonstrate the generability of our method, we evaluate the performance of SVD-LLM and the baselines on seven models from three different LLM families (LLaMA-7B, 13B, 30B, 65B (Touvron et al., 2023a), LLaMA2-7B (Touvron et al., 2023b), Vicuna-7B (Chiang et al., 2023) and Mistral-7B (Jiang et al., 2023) and $11$ datasets including three language modeling datasets (WikiTex-2 (Merity et al., 2016), PTB (Marcus et al., 1993) and C4 (Raffel et al., 2020)) and eight classification datasets (SST-5 (Socher et al., 2013), MNLI (Williams et al., 2018), MRPC (Dolan & Brockett, 2005), COLA (Warstadt et al., 2019), SNLI (Bowman et al., 2015), ARC-e, ARC-c, and ARC-b (Clark et al., 2018)) in zero-shot setting following previous experimental design (Ma et al., 2023).
@@ 4.2 Overall Performance
LoRA (Hu et al., 2021) is a common fine-tuning technique for LLM. It has been applied with pruning-based LLM compression methods such as LLM-Pruner (Ma et al., 2023) to mitigate accuracy drop after pruning. LoRA can also be combined with SVD-based LLM compression methods by modifying the forward pass of a linear layer as:
@@ 4.5 Benefits to other LLM Compression Methods
SVD-LLM is orthogonal to other LLM compression methods including quantization and parameter pruning. In this experiment, we add the techniques proposed in SVD-LLM into state-of-the-art quantization and parameter pruning-based LLM compression methods to examine how SVD-LLM could further enhance their performance.
@@
Method GPTQ-4bit GPTQ-3bit SVD-LLM + GPTQ-4bit
@@
Method LLM-Pruner-30% LLM-Pruner-40%

### 15 Mar 2024 v2

@@ 1 Introduction
We compare SVD-LLM with three SVD-based methods for LLM compression, including vanilla SVD as well as state-of-the-art methods FWSVD and ASVD. To demonstrate the generability of SVD-LLM, we conduct our evaluation on a total of $11$ datasets and seven models from three different LLM families (LLaMA, Vicuna, and Mistral) at four different scales (7B, 13B, 30B, and 65B). We highlight five of our findings: (1) SVD-LLM consistently outperforms vanilla SVD, FWSVD, and ASVD across all $11$ datasets, three different LLM families, and four different scales. (2) None of the existing SVD-based LLM compression methods report their performance on model compression ratios above 20%. SVD-LLM exhibits significant advantages over vanilla SVD, FWSVD, and ASVD under high compression ratios from 30% to 60%. (3) SVD-LLM also exhibits superiority over state-of-the-art in terms of compression speed. Specifically, when compressing LLaMA-7B under 20% compression ratio on an A100 GPU, ASVD takes about $5.5$ hours whereas SVD-LLM completes the compression process in $15$ minutes. (4) Moreover, SVD-LLM can benefit other LLM compression methods. Our evaluation results show that SVD-LLM is able to further enhance the compression performance of well-recognized quantization (GPTQ (Frantar et al., 2022)) and parameter pruning-based (LLM-Pruner (Ma et al., 2023)) LLM compression methods. (5) Lastly, SVD-LLM brings additional benefit beyond compressing the sizes of LLMs, and is also able to reduce the footprint of KV cache during inference at runtime.
@@ 3.1 Truncation-Aware Data Whitening
Key Design: The key idea of SVD-LLM is to incorporate a truncation-aware data whitening technique that ensures a direct mapping between singular values and compression loss. To achieve this, SVD-LLM change the whitened activation $\tilde{X}=S^{-1}X$ to be orthogonal so that each channel is independent of each other, i.e., $\tilde{X}\tilde{X}^{T}=S^{-1}XX^{T}(S^{-1})^{T}=I$. Though the matrix $S$ is not unique, we found that the Cholesky decomposition (Meyer, 2000) of $XX^{T}$ is a good solution due to its numerical stability. After that we perform SVD on $\tilde{W}=WS$ instead of the original weight $W$ to obtain three decomposed matrices $\tilde{U},\tilde{\Sigma},\tilde{V}$, where $\tilde{U}=[\tilde{u_{1}},\tilde{u_{2}},\tilde{u_{3}},...,\tilde{u_{r}}]$, $\tilde{\Sigma}=\text{diag}(\tilde{\sigma_{1}},\tilde{\sigma_{2}},\tilde{\sigma_{3}},\cdots,\tilde{\sigma_{r}})$, and $\tilde{V}=[\tilde{v_{1}},\tilde{v_{2}},\tilde{v_{3}},...,\tilde{v_{r}}]$. Lastly, the smallest singular values are truncated to compress $\tilde{\Sigma}$ to $\tilde{\Sigma}_{1}$ and obtain the compressed weight $\tilde{W^{\prime}}=\tilde{U^{\prime}}\tilde{\Sigma}_{1}\tilde{V}^{T}$.
@@ 4.1 Experimental Setups
Models and Datasets. To demonstrate the generability of our method, we evaluate the performance of SVD-LLM and the baselines on seven models from three different LLM families (LLaMA-7B, 13B, 30B, 65B (Touvron et al., 2023a), LLaMA2-7B (Touvron et al., 2023b), Vicuna-7B (Chiang et al., 2023) and Mistral-7B (Jiang et al., 2023) and $11$ datasets including three language modeling datasets (WikiTex-2 (Merity et al., 2017), PTB (Marcus et al., 1993) and C4 (Raffel et al., 2020)) and eight classification datasets (SST-5 (Socher et al., 2013), MNLI (Williams et al., 2018), MRPC (Dolan & Brockett, 2005), COLA (Warstadt et al., 2019), SNLI (Bowman et al., 2015), ARC-e, ARC-c, and ARC-b (Clark et al., 2018)) in zero-shot setting following previous experimental design (Ma et al., 2023).
@@ 4.2 Overall Performance
LoRA (Hu et al., 2022) is a common fine-tuning technique for LLM. It has been applied with pruning-based LLM compression methods such as LLM-Pruner (Ma et al., 2023) to mitigate accuracy drop after pruning. LoRA can also be combined with SVD-based LLM compression methods by modifying the forward pass of a linear layer as:
@@ 4.5 Benefits to other LLM Compression Methods
SVD-LLM is orthogonal to other LLM compression methods including quantization and parameter pruning. In this experiment, we add the techniques proposed in SVD-LLM into representative quantization and parameter pruning-based LLM compression methods that are widely recognized by the community to examine how SVD-LLM could further enhance their performance.
@@
Metric GPTQ-4bit GPTQ-3bit SVD-LLM + GPTQ-4bit
@@
Metric LLM-Pruner-30% LLM-Pruner-40%